OpenStudy (anonymous):

Solve for x. (round to 3 decimal places). e^x=5

OpenStudy (anonymous):

HINT: Add ln to both sides.

OpenStudy (anonymous):

\[\large{\ln e^x = \ln 5}\]

OpenStudy (anonymous):

hmmm, so would I do something like ln+e^5=ln+5?

OpenStudy (anonymous):

i got answer of 1.609 is whatg should = to... Does that sound right?

OpenStudy (anonymous):

Actually I mean take the natural logarithm of both sides. Sorry for confusion.

OpenStudy (anonymous):

Yup, this is correct.

OpenStudy (anonymous):

x = 1.609

OpenStudy (anonymous):

awesome, thank you for you help and fast responses :)

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

makes sense now, lol